A α  alpha
B β  beta
Γ γ  gamma
Δ δ  delta
E ϵ, ε  epsilon
Z ζ  zeta
H η  eta
Θ θ, ϑ  theta
I ι iota
K κ  kappa
Λ λ  lambda
M μ  mu
N ν  nu
Ξ ξ  xi
O o  omikron
Π π, ϖ  pi
P ρ, ϱ  rho
Σ σ, ς  sigma
T τ  tau
Υ υ upsilon, ypsilon
Φ ϕ, φ  phi
X χ  chi
Ψ ψ  psi
Ω ω  omega